byeeee:

Given the function g(x) = 6(4)x, Section A is from x = 0 to x = 1 and Section B is from x = 2 to x = 3. Part A: Find the average rate of change of each section. (4 points) Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other. (6 points)

byeeee:

@dude

dude:

Do \(g(x) = 6(4)^0\) \(g(x) = 6(4)^1\) \(g(x) = 6(4)^2\) \(g(x) = 6(4)^3\)